# Era3D: High-Resolution Multiview Diffusion using Efficient Row-wise Attention

### v2

@@ Abstract
In this paper, we introduce Era3D, a novel multiview diffusion method that generates high-resolution multiview images from a single-view image. Despite significant advancements in multiview generation, existing methods still suffer from camera prior mismatch, inefficacy, and low resolution, resulting in poor-quality multiview images. Specifically, these methods assume that the input images should comply with a predefined camera type, e.g. a perspective camera with a fixed focal length, leading to distorted shapes when the assumption fails. Moreover, the full-image or dense multiview attention they employ leads to an exponential explosion of computational complexity as image resolution increases, resulting in prohibitively expensive training costs. To bridge the gap between assumption and reality, Era3D first proposes a diffusion-based camera prediction module to estimate the focal length and elevation of the input image, which allows our method to generate images without shape distortions. Furthermore, a simple but efficient attention layer, named row-wise attention, is used to enforce epipolar priors in the multiview diffusion, facilitating efficient cross-view information fusion. Consequently, compared with state-of-the-art methods, Era3D generates high-quality multiview images with up to a 512$\times$512 resolution while reducing computation complexity by 12x times. Comprehensive experiments demonstrate that Era3D can reconstruct high-quality and detailed 3D meshes from diverse single-view input images, significantly outperforming baseline multiview diffusion methods. Project page: https://penghtyx.github.io/Era3D/.
@@ 1 Introduction
3D reconstruction from single-view images is an essential task in computer vision and graphics due to its potential applications in game design, virtual reality, and robotics. Early research [68, 70, 79, 118] mainly relies on direct 3D regression on voxels [64, 88, 16], which often leads to oversmoothed results and has difficulty in generalizing to real-world unseen objects due to limited 3D training data [7]. Recently, diffusion models (DMs) [27, 75] show strong generation ability on image or video synthesis by training on extremely large-scale datasets. These diffusion models are promising tools for single-view 3D reconstruction because it is possible to generate novel-view images from the given image to enable 3D reconstruction.
@@
Despite impressive advancements in multiview diffusion methods [84, 55, 53, 80, 52, 51, 103, 95, 95, 94, 112, 45], efficiently generating novel-view images for high-quality 3D reconstruction remains an open challenge. There are three noticeable challenges in the current multiview diffusion methods. (1) Inconsistent predefined camera type. Most multiview diffusion methods assume that the input images are captured by a camera with a predefined focal length. This leads to unwanted distortions when input images are captured by cameras with different camera types or intrinsics, as exemplified in Fig. 3 (e.g., Wonder3D’s assumption of an orthogonal camera leads to distorted meshes when the input image is captured by a perspective camera with a small focal length). (2) Inefficiency of multiview diffusion. Multiview diffusion methods usually rely on multiview attention layers to exchange information among different views to generate multiview-consistent images. However, these multiview attention layers are usually implemented by extending the self-attention in Stable Diffusion [78] to all multiview images, which is called dense multiview attention in Fig.2(a) and results in a significant increase in computation complexity and memory consumption. (3) Low resolution of generated images. The limitation above restricts most existing multiview diffusion models to resolutions of $256\times 256$, preventing them from reconstructing detailed meshes. Addressing these above challenges is crucial for developing practical and scalable multiview diffusion methods.
@@
In this paper, we introduce Era3D, a novel multiview diffusion method that efficiently generates high-resolution (512$\times$512) consistent-multiview images for single-view 3D reconstruction. Unlike existing methods, Era3D allows images of commonly used camera types as inputs while mitigating the unwanted distortion brought by different camera models. To this end, we employ a unique approach: using different camera models for input images and the generated ones for training, meaning that the input images are allowed to have arbitrary focal lengths and elevations while generated images are with orthogonal cameras and fixed viewpoints of 0 elevations.
         Input   Generated images and mesh
@@
However, this requires DMs to implicitly infer and rectify the focal lengths and viewpoints of input images in the generation process, which is a challenging task and degenerates the generation quality. To overcome this challenge and improve generation quality, we propose a novel regression and condition scheme and utilize the low-level feature maps of UNet at each denoising step to predict camera information of input images. We find that such a regression and condition scheme facilitates much more accurate camera pose prediction than existing methods [51] and leads to more details in the generation. As shown in Fig. 3, Era3D successfully avoids the above distortion problem brought by the different camera types and focal lengths.
@@
Moreover, drawing inspiration from epipolar attention [98], Era3D enables efficient training for high-resolution multiview generation by introducing a novel row-wise multiview attention. Epipolar constraint can be utilized to constrain the attention regions across views and thus improve attention efficiency. However, directly applying such epipolar attention [98] for a general camera setting (Fig. 2(b)) is still memory and computationally inefficient because we have to sample multiple points on epipolar lines for attention. This requires us to construct a 3D grid of features in the view frustums for multiview images, which is too slow and memory-consuming. In contrast, since Era3D generates images with orthogonal cameras on viewpoints of $0^{\circ}$, we find that epipolar lines in our camera setting are aligned with rows of images across different views (Fig. 2(d)), which enables us to propose an efficient row-wise attention layer. Compared with dense multiview attention, row-wise attention significantly reduces memory consumption ($35.32$GB v.s. $1.66$GB), and the computation complexity ($220.41$ms v.s. $2.23$ms) of multiview attention (Fig. 2(e)). Even with Xformers [44], an accelerating library for attention, the efficiency of row-wise attention still outperforms existing methods by approximately twelve-fold as evident in Tab. 3. Consequently, the proposed row-wise attention allows us to easily scale Era3D to a high resolution of 512$\times$512 to reconstruct more detailed 3D meshes.
@@
Overall, our main contributions are summarized as follows: (1) Era3D is the first method that tries to solve the distortion artifacts brought by the inconsistent camera intrinsic in 3D generation; (2) we design a novel regression and condition scheme to enable diffusion models to take images of arbitrary cameras as inputs while outputting the orthogonal images on the canonical camera setting; (3) we propose row-wise multiview attention, an efficient attention layer for high-resolution multiview image generation; (4) our method achieve state-of-the-art performance for single-view 3D generation [18].
@@ 2 Related Works
Our study is primarily centered on the domain of image-to-3D. Unlike early works [10, 73, 92, 54, 106, 81, 82, 48, 97, 31, 13, 80, 32, 25, 2, 37, 105, 14, 113, 46, 17, 47, 49, 74, 114], which concentrate on per-scene optimization based on Score Distillation Sampling [72, 100], we emphasize feed-forward 3D generation.
Image to 3D. Generating 3D assets from images has been extensively researched, paralleling the development of GAN [23, 38, 40, 41, 39, 61, 63, 19, 86, 5, 22, 35] and diffusion models (DMs) [87, 28, 34, 62, 65, 101, 20, 76, 29]. A stream of these works [33, 119, 8, 30, 112], directly produce 3D representations, like SDF [69, 16, 67], NeRF [59, 24, 115, 9, 56, 116, 3], Triplane [6, 26, 85], Gaussian [42, 89, 1, 109] or 3D volume [90]. Zero-1-to-3 [52] and subsequent works [83, 51, 12] represent the scene as a diffusion model conditioned on reference image and camera pose. LRM-based mthods [30, 45, 104, 110, 111, 112] employ a large transformer architecture to train a triplane representation with a data-driven approach. Another technical line involves generating consistent multiview images first [84, 103, 83, 93, 96, 107], and then robustly reconstructing 3D shapes with NeuS [53, 55], Gaussian Splatting [91, 30, 111] or LRM [108]. Despite great advancement, challenges remain in reconstruction quality, training resolution, or efficiency.
Multiview Diffusion. Cross-view consistency is critical in 3D reconstruction and generation, relying on multiview feature correspondence to estimate 3D structures. MVffusion [95] first proposes generating multiview images in parallel with correspondence-aware attention, facilitating cross-view information interaction, and applies to texture scene meshes. [98, 36] introduces epipolar features into DM to enhance fusion between viewpoints. Zero123++ [83] tiles multi-views into a single image and performs a single pass for multiview generation, also used in Direct2.5 and Instant3D. MVDream [84] and Wonder3D [55] also design multiview self-attention to improve multiview consistency. Syncdreamer [53] composes multiview features into 3D volumes, conducting 3D-aware fusion in 3D noise space. All of the aforementioned methods share the same idea: modeling 3D generation with multiview joint probability distribution. Other works [15, 57, 99] explore the priors from video diffusion model to achieve consistent multiview generation. Following widely used multiview self-attention, we propose more efficient row-wise attention to reduce computation workloads, but without loss of multiview feature interaction.
@@ 3.2 Regression and Condition Scheme
Given an image with an arbitrary viewpoint and focal length, generating novel-view images in the canonical camera setting is challenging because this implicitly puts an additional task on the diffusion model to infer the focal lengths and elevation of the input image. To make it easier, previous methods [52, 53, 4] rely on additional elevation inputs or predictions as the conditions for the diffusion model. However, pose estimation from a single image is inherently ill-posed due to the lack of geometry information. Even though estimating a rough elevation is possible, it is almost impossible for users to estimate the focal length of the input image.
@@ 4 Experiments
Implementation details. Our implementation is built upon the open-source text-to-image model, SD2.1-unclip [77]. We train Era3D on $16$ H800 GPUs (each with $80$ GB) using a batch size of $128$ for $40,000$ steps. We set the initial learning rate as 1e-4 and decreased it to 5e-5 after $5,000$ steps. The training process takes approximately $30$ hours. To conduct classifier-free guidance (CFG) [29], we randomly omit the clip condition at a rate of $0.05$. During inference, we employ the DDIM sampler [87] with $40$ steps and a CFG scale of $3.0$ for the generation. Following Wonder3D [55], we first perform reconstruction with NeuS and then apply a texture refinement step. For further details, please refer to the supplementary material.
@@ 4.1 Experimental Results
RealFusion 0.0819 0.2714 0.283
Zero-1-to-3 0.0339 0.5035 0.166
One-2-3-45 0.0629 0.4086 -
Shap-E 0.0436 0.3584 -
Magic123 0.0516 0.4528 -
SyncDreamer 0.0261 0.5421 0.146
Wonder3D 0.0248 0.5678 0.141
LGM 0.0259 0.5628 -
Ours 0.0217 0.5973 0.126
@@ A.1 Implementation Details
Reconstruction and texture refinement. We perform reconstruction from the generated multiview normal and color images using NeuS. However, since we only generate sparse views, the rendered images of the reconstructed meshes may not attain the quality of the generated ones, particularly for objects with complex textures. To address this issue, we employ differential rendering to refine the textures on the reconstructed meshes. Specifically, we preserve the geometry of Neus and initialize the texture with vertex colors. We utilize [43] to render multiview images in predefined views and optimize the vertex colors with the corresponding generated images. This process takes less than $1$ second and significantly improves the texture quality.
@@ A.3 Pose Estimation
$f$=35 $f$=50 $f$=85 $f$=105 $f$=135 $\alpha$=-10 $\alpha$=10 $\alpha$=20 $f$=30 $f$=40

### v3

@@ Abstract
In this paper, we introduce Era3D, a novel multiview diffusion method that generates high-resolution multiview images from a single-view image. Despite significant advancements in multiview generation, existing methods still suffer from camera prior mismatch, inefficacy, and low resolution, resulting in poor-quality multiview images. Specifically, these methods assume that the input images should comply with a predefined camera type, e.g. a perspective camera with a fixed focal length, leading to distorted shapes when the assumption fails. Moreover, the full-image or dense multiview attention they employ leads to a dramatic explosion of computational complexity as image resolution increases, resulting in prohibitively expensive training costs. To bridge the gap between assumption and reality, Era3D first proposes a diffusion-based camera prediction module to estimate the focal length and elevation of the input image, which allows our method to generate images without shape distortions. Furthermore, a simple but efficient attention layer, named row-wise attention, is used to enforce epipolar priors in the multiview diffusion, facilitating efficient cross-view information fusion. Consequently, compared with state-of-the-art methods, Era3D generates high-quality multiview images with up to a 512$\times$512 resolution while reducing computation complexity of multiview attention by 12x times. Comprehensive experiments demonstrate the superior generation power of Era3D- it can reconstruct high-quality and detailed 3D meshes from diverse single-view input images, significantly outperforming baseline multiview diffusion methods. Project page: https://penghtyx.github.io/Era3D/.
@@ 1 Introduction
3D reconstruction from single-view images is an essential task in computer vision and graphics due to its potential applications in game design, virtual reality, and robotics. Early research [42, 44, 53, 78, 28] mainly relies on direct 3D regression on voxels [39, 58, 9], which often leads to oversmoothed results and has difficulty in generalizing to real-world unseen objects due to limited 3D training data [4]. Recently, diffusion models (DMs) [16, 50] show strong generation ability on image or video synthesis by training on extremely large-scale datasets [47, 48]. These diffusion models are promising tools for single-view 3D reconstruction because it is possible to generate novel-view images from the given image to enable 3D reconstruction.
@@
         Input   Generated images and mesh
Despite impressive advancements in multiview diffusion methods [55, 34, 33, 32, 31, 64, 63, 27], efficiently generating novel-view images for high-quality 3D reconstruction remains an open challenge. There are three noticeable challenges in the current approaches. (1) Inconsistent predefined camera type. Most multiview diffusion methods assume that the input images are captured by a camera with a predefined focal length. This leads to unwanted distortions when input images are captured by cameras with different camera types or intrinsics, as exemplified in Fig. 2 (e.g., Wonder3D’s assumption of an orthogonal camera leads to distorted meshes when the input image is captured by a perspective camera with a small focal length). (2) Inefficiency of multiview diffusion. Multiview diffusion methods usually rely on multiview attention layers to exchange information among different views to generate multiview-consistent images. However, these multiview attention layers are usually implemented by extending the self-attention in Stable Diffusion [52] to all multiview images, which is called dense multiview attention in Fig.3(a) and results in a significant increase in computation complexity and memory consumption. (3) Low resolution of generated images. The limitation above restricts most existing multiview diffusion models to resolutions of $256\times 256$, preventing them from reconstructing detailed meshes. Addressing above challenges is crucial for developing practical and scalable multiview diffusion methods.
@@
In this paper, we introduce Era3D, a novel multiview diffusion method that efficiently generates high-resolution (512$\times$512) consistent-multiview images for single-view 3D reconstruction. Unlike existing methods, Era3D allows images of commonly used camera types as inputs while mitigating the unwanted distortion brought by different camera models.
@@
To this end, we employ a unique approach: using different camera models for input images and the generated ones for training, meaning that the input images are allowed to have arbitrary focal lengths and elevations while generated images are with orthogonal cameras and fixed viewpoints of 0 elevations. However, this requires DMs to implicitly infer and rectify the focal lengths and viewpoints of input images in the generation process, which is a challenging task and degrades the generation quality. To overcome this challenge and improve generation quality, we propose a novel regression and condition scheme and utilize the low-level feature maps of UNet at each denoising step to predict camera information of input images. We find that such a regression and condition scheme facilitates much more accurate camera pose prediction than existing methods [31] and leads to more details in the generation. As shown in Fig. 2, Era3D successfully avoids the above distortion problem brought by the different camera types and focal lengths.
@@
Moreover, drawing inspiration from epipolar attention [65], Era3D enables efficient training for high-resolution multiview generation by introducing a novel row-wise multiview attention. Epipolar constraint can be utilized to constrain the attention regions across views and thus improve attention efficiency. However, directly applying such epipolar attention [65] for a general camera setting (Fig. 3(b)) is still memory and computationally inefficient because one would have to sample multiple points on epipolar lines for attention. This would require to construct a 3D grid of features in the view frustums for multiview images, which is too slow and memory-consuming. In contrast, since Era3D generates images with orthogonal cameras on viewpoints of $0^{\circ}$, we find that epipolar lines in our camera setting are aligned with pixel rows of images across different views (Fig. 3(d)), which enables us to propose an efficient row-wise attention layer. Compared with dense multiview attention, row-wise attention significantly reduces memory consumption ($35.32$GB v.s. $1.66$GB), and the computation complexity ($220.41$ms v.s. $2.23$ms) of multiview attention (Fig. 3(e)). Even with Xformers [26], an accelerating library for attention, the efficiency of row-wise attention still outperforms existing methods by approximately twelve-fold as evident in Tab. 3. Consequently, the proposed row-wise attention allows us to easily scale Era3D to a high resolution of 512$\times$512 to reconstruct more detailed 3D meshes.
@@
Overall, our main contributions are summarized as follows: (1) Era3D is the first method that tries to solve the distortion artifacts brought by the inconsistent camera intrinsic in 3D generation; (2) we design a novel regression and condition scheme to enable diffusion models to take images of arbitrary cameras as inputs while outputting the orthogonal images on the canonical camera setting; (3) we propose row-wise multiview attention, an efficient attention layer for high-resolution multiview image generation; (4) our method achieves state-of-the-art performance for single-view 3D generation [10].
@@ 2 Related Works
Our study is primarily centered on the domain of image-to-3D. Unlike early works [6, 49, 62, 71, 29], which concentrate on per-scene optimization based on Score Distillation Sampling [46, 67], we emphasize feed-forward 3D generation.
Image to 3D. Generating 3D assets from images has been extensively researched, paralleling the development of GAN [13, 23, 38] and diffusion models (DMs) [57, 17, 11, 18]. A stream of these works [20, 79, 5, 19, 76], directly produce 3D representations, like SDF [43, 9, 41], NeRF [36, 14, 1], Triplane [3, 15, 56], Gaussian [24, 59] or 3D volume [60]. Zero-1-to-3 [32] and subsequent works [54, 31] represent the scene as a diffusion model conditioned on reference image and camera pose. LRM-based methods [19, 27, 70, 74] employ a large transformer architecture to train a triplane representation with a data-driven approach. Another technical line involves generating consistent multiview images first [55, 69, 54, 72], and then robustly reconstructing 3D shapes with NeuS [33, 34], Gaussian Splatting [61, 19, 75] or LRM [73]. Despite significant advancements, challenges remain in reconstruction quality, training resolution, or efficiency.
Multiview diffusion. Cross-view consistency is critical in 3D reconstruction and generation, relying on multiview feature correspondence to estimate 3D structures. MVffusion [64] first proposes generating multiview images in parallel with correspondence-aware attention, facilitating cross-view information interaction, and applies to texture scene meshes. [65, 22] introduces epipolar features into DM to enhance fusion between viewpoints. Zero123++ [54] tiles multi-views into a single image and performs a single pass for multiview generation, which is also used in Direct2.5 and Instant3D. MVDream [55] and Wonder3D [34] also design multiview self-attention to improve multiview consistency. Syncdreamer [33] composes multiview features into 3D volumes, conducting 3D-aware fusion in 3D noise space. All of the aforementioned methods share the same idea: modeling 3D generation with multiview joint probability distribution. Other works [8, 66] explore the priors from video diffusion model to achieve consistent multiview generation. Following widely used multiview self-attention, we propose more efficient row-wise attention to reduce computation workloads, but without loss of multiview feature interaction.
@@ 3.2 Regression and Condition Scheme
Given an image with an arbitrary viewpoint and focal length, generating novel-view images in the canonical camera setting is challenging because this implicitly puts an additional task on the diffusion model to infer the focal lengths and elevation of the input image. To make it easier, previous methods [32, 33, 2] rely on additional elevation inputs or predictions as the conditions for the diffusion model. However, pose estimation from a single image is inherently ill-posed due to the lack of geometry information. Moreover, even though estimating a rough elevation is possible, it is almost impossible for users to estimate the focal length of the input image.
@@ 4 Experiments
Implementation details. Our implementation is built upon the open-source text-to-image model, SD2.1-unclip [51]. We train Era3D on $16$ H800 GPUs (each with $80$ GB) using a batch size of $128$ for $40,000$ steps. We set the initial learning rate as 1e-4 and decreased it to 5e-5 after $5,000$ steps. The training process takes approximately $30$ hours. To conduct classifier-free guidance (CFG) [18], we randomly omit the clip condition at a rate of $0.05$. During inference, we employ the DDIM sampler [57] with $40$ steps and a CFG scale of $3.0$ for the generation. Following Wonder3D [34], we first perform reconstruction with NeuS and then apply a texture refinement step. The whole pipeline requires approximately 4 minutes, comprising 13 seconds for the multiview diffusion, 3 minutes for the NeuS reconstruction, and 10 seconds for the texture refinement. For further details, please refer to the supplementary material.
@@ 4.1 Experimental Results
           GSO                  In-the-wild
   Input      Ours    Wonder3D     Unique3D       Input      Ours     Wonder3D     Unique3D
Figure 7: More comparisons w.r.t distortion problem.
Distortion problem We finally provide the comparisons with SOTA methods, like Unique3D, w.r.t distortion problem in Fig. 7, which shows that they suffers from severe perspective distortion while our method greatly alleviates this problem.
@@
RealFusion 0.0819 0.2714 0.283 0.722 15.26
Zero-1-to-3 0.0339 0.5035 0.166 0.779 18.93
One-2-3-45 0.0629 0.4086 - - -
Shap-E 0.0436 0.3584 - - -
Magic123 0.0516 0.4528 - - -
SyncDreamer 0.0261 0.5421 0.146 0.798 20.05
Wonder3D 0.0248 0.5678 0.141 0.811 20.83
LGM 0.0259 0.5628 - - -
Ours 0.0217 0.5973 0.126 0.837 22.74
@@ 4.3 Ablations and Discussions
Finally, we conduct performance comparisons between dense, epipolar, and our row-wise attention. Considering that in our orthogonal setup with the same elevation, epipolar attention is equivalent to row-wise attention (except for implementation differences), we compared dense and row-wise attention at a resolution of 256. Era3D achieves comparable performance with dense multiview attention, as evidenced in Tab. 4. Our row-wise setting outperforms the baseline for the Chamfer distance, LPIPS, and PSNR. We attribute this to row-wise attention reducing the number of attention tokens, allowing the model to focus more on valuable tokens.
@@
Table 4: Performance of dense and row-wise attention at a resolution of 256.
Method CD ↓ IoU ↑ LPIPS ↓ PSNR ↑ SSIM ↑
dense 0.0239 0.5877 0.140 20.73 0.819
rowwise 0.0232 0.5831 0.137 20.92 0.813
@@ 5 Limitation and Conclusion
6 Ethics Statement
The objective of Era3D is to equip users with a powerful tool for creating detailed 3D models. Our method allows users to generate 3D objects based on a single image. However, there is a potential risk that these generated models could be misused to deceive viewers. It is important to note that this issue is not unique to our methodology but prevalent in other generative model methodologies. Therefore, it is absolutely essential for current and future research in the field of 3D generative modeling to address and reassess these considerations consistently.
Acknowledgments and Disclosure of Funding
The research was supported by Theme-based Research Scheme (T45-205/21-N) from Hong Kong RGC, and Generative AI Research and Development Centre from InnoHK.
@@ A.1 Implementation Details
Reconstruction and texture refinement. We perform reconstruction from the generated multiview normal and color images using NeuS. However, since we only generate sparse views, the rendered images of the reconstructed meshes may not attain the quality of the generated ones, particularly for objects with complex textures. To address this issue, we employ differential rendering to refine the textures on the reconstructed meshes. Specifically, we preserve the geometry of Neus and initialize the texture with vertex colors. We utilize [25] to render multiview images in predefined views and optimize the vertex colors with the corresponding generated images. This process takes less than $1$ second and significantly improves the texture quality. For the NeuS reconstruction, we use the same settings as Wonder3D. During texture refinement, we optimize the appearance for 200 iterations with a resolution of 1024 and a learning rate of 1e-3.
@@ A.3 Pose Estimation
$f$=35 $f$=50 $f$=85 $f$=105 $f$=135 $\alpha$=-10 $\alpha$=10 $\alpha$=20 $\alpha$=30 $\alpha$=40